# SReferences

### v1

@@ 5.4 Ablation Studies
Tab. 2 displays the results of ablation studies. In the first ablation study, we applied our idempotent strategy (with identity decoder) using the EfficientNetAutoAttB4ST features. In the second study, we concatenated the features from the identity-aware feature extractor with the features of EfficientNetAutoAttB4ST as we did in our proposed method and used a feedforward network to classify the images. The first ablation achieved the AUC mean of $0.926$ and the AUC median was $0.928$. The sample standard deviation and interquartile range were $0.05$ and $0.06$. The second ablation achieved the AUC mean of $0.893$ and the AUC median was $0.920$. The sample standard deviation and interquartile range were $0.10$ and $0.13$. The achieved AUC values are much lower compared to the proposed method. This confirms that the identity conditioning and idempotence strategy have synergy (positive interaction).
Table 2: Ablation studies for the proposed method.
@@
6 Discussion
In the current work, the reconstruction model is the deepfake generation tool, i.e., Faceswap-GAN, one of the most popular and effective off-the-shelf tools for deepfake generation. When the reconstruction model does not match the tool used for deepfake generation, the processing traces caused by deepfake and reconstruction may be different, but both traces are generated by neural networks. For example, the reconstruction model is a Faceswap-GAN model whereas the deepfake video is generated by a Faceswap model. A more sophisticated classifier may be needed to exploit the processing traces left by different network structures. Under the proposed double operations framework, a more general Siamese neural network for processing-trace manifold learning may work, but we leave this exploration to future work.
@@ 6 Discussion
Compared to end-to-end CNN-based classifiers, our proposed method targets deepfake detection for individuals, with main applications on public figures. Although our method needs training the reconstruction models, the training can be done in advance for each public figure. For example, a journalist can train the reconstruction models for various candidates before they need to verify videos for reporting tasks. Journalists may also share or collaboratively train detectors within their professional networks. To let the detection system support a new individual, the journalist will need to train a reconstruction operator for that individual and then finetune the Siamese network.
7 Conclusion and Future Work
@@
In this work, we have proposed to use the method of double neural network operations and individual conditioning for the deepfake detection. The proposed detector can achieve better detection performance than end-to-end CNN-based detectors on our curated dataset of public figures with identity labels. We have found that utilizing identity information can make the deepfake detector more reliable. In future work, we plan to extend the double-operations detection to scenarios with mismatched neural network architectures.
Supplementary Material
1 Tool for Deepfake Generation
[…]
2 Deepfake Generation and Reconstruction Model $R$
For each target video, we generated a deepfake video by feeding a Faceswap-GAN model with the target video and a video of a public figure with a known identity. The video of the public figure was collected from YouTube and has the same gender as in the template video. For deepfake generation, we swap the face of the public figure onto the face of the person in the target video since public figures are usually the victims of deepfakes.
For deepfake detection using double operations, a journalist does not know which videos were used to generate the potentially fake video and knows only the identity of the video in question. Therefore, the reconstruction model was trained using videos of the known public figure from scenes other than the videos used for deepfake generation. The reconstructing unit $R$ is chosen to share the same neural network structure as the deepfake generating unit, i.e., a Faceswap-GAN. For a specific public figure, we trained a Faceswap-GAN using five videos, with $40,\!000$ iterations to ensure a good reconstruction quality.
Figure 6: A schematic for autoencoder-based deepfake generator. The two encoders have shared weights and are used to extract features of the faces from both persons A and B. The decoders for A and B are trained to reconstruct the faces of persons A and B, respectively. A dashed arrow represents the deepfake generation step, i.e., feature A is sent to decoder B and the reconstructed image is the face of person B with facial expressions from person A.
3 Advantage of Identity-Conditioned Feature Extraction
@@
[…]
@@
[…]
@@
[…]
@@
4 Fine-Grained Performance Analysis Over Identities
@@
[…]

### v2

@@ V-E Ablation Studies
TABLE IV displays the results of ablation studies. In the first ablation study, we applied our idempotent strategy (with identity decoder) using the EfficientNetAutoAttB4ST features. In the second study, we concatenated the features from the identity-aware feature extractor with the features of EfficientNetAutoAttB4ST as we did in our proposed method and used a feedforward network to classify the images. The first ablation achieved the AUC mean of $0.926$, and the AUC median was $0.928$. The sample standard deviation and interquartile range were $0.05$ and $0.06$. The second ablation achieved the AUC mean of $0.893$, and the AUC median was $0.920$. The sample standard deviation and interquartile range were $0.10$ and $0.13$. The achieved AUC values are much lower compared to the proposed method. This confirms that the identity conditioning and idempotence strategy have synergy (positive interaction).
TABLE IV: Ablation studies for the proposed method.
@@
V-F Experimental Verification of Near-Idempotence
Our proposed detection method leverages the near-idempotence property of the deepfake operator. Exact idempotence occurs when an altered image, passed through a deepfake generator, depicts no further changes. In the case of near-idempotence, the second operation would lead to small changes compared to the first operation. Let the residues be defined concerning raw data $f$ as follows:
$e_{0}$ $=R_{\text{recon}}(f)-f,$ (5a)
$e_{1}$ $=R_{\text{recon}}(R_{\text{df}}(f))-R_{\text{df}}(f).$ (5b)
To establish near-idempotence, we require $\|e_{0}\|_{2}\gg\|e_{1}\|_{2}$ for all $f$, $R_{\text{df}}$, and $R_{\text{recon}}$, where $R_{\text{df}}$ represents the deepfake operation and $R_{\text{recon}}$ represents the reconstruction operation. In this subsection, we experimentally verify this property of deepfake generators. Specifically, we focus on two types of deepfake operations: Faceswap-GAN (FG) and diffusion-based (D) methods. Based on the choice and the order of deepfake operations applied to an image, we present our results within three categories as follows.
V-F1 $R_{\mathrm{df}}=R_{\mathrm{recon}}=FG$
We trained a Faceswap-GAN face reconstructor for each identity using a subset of the available faces of that identity from the CelebDF dataset. Using these reconstructors for both deepfake generation and double neural network operation, we illustrate the residual vectors in Fig. 6 (a) and Fig. 6 (b). The residual vectors $e_{0}$ corresponding to an authentic image processed by an operator are shown in Fig. 6 (a). When the singly processed image is processed again by the same operator, the residual vectors $e_{1}$ are shown in Fig. 6 (b).
(a) Refer to caption Refer to caption (b)                                      (c)
Figure 6: t-SNE visualization for residual vectors (a) $e_{0}=\mathrm{FG}(f)-f$ and (b) $e_{1}=\mathrm{FG}(\mathrm{FG}(f))-\mathrm{FG}(f)$, when is Faceswap-GAN (FG) is the deepfake operation in question. We randomly highlighted only 5% vectors in opaque color for better viewing experiences. Contrasting (a) with (b), we note that residual vectors are shorter when the same operator is applied twice. The same can be concluded by contrasting residual vectors (a) $e_{0}=\mathrm{FG}(f)-f$ and (c) $e_{1}=\mathrm{FG}(\mathrm{D}(f))-\mathrm{D}(f)$ when diffusion (D) is the deepfake operation in question.
These plots reveal that the first operation results in significant vector shifts, whereas the second operation leads to minimal shifts for $81.3\%$ of the samples. While Fig. 6 presents results for the CACD dataset, the corresponding plots for the CelebDF dataset are provided in the supplementary document.
V-F2 $R_{\mathrm{df}}=R_{\mathrm{recon}}={D}$
To verify the near-idempotence property of diffusion-based deepfake generators, we applied two consecutive diffusion operations to authentic images. In this case, the repeated operation yields small residual vectors for $89.1\%$ of the samples. The residual vectors are demonstrated in the supplementary document.
V-F3 $R_{\mathrm{df}}={D}\,\,\&\,\,R_{\mathrm{recon}}={FG}$
In this case, the operations are diffusion followed by Faceswap-GAN. The residual vectors for the diffusion-generated deepfakes are shown in Fig. 6 (c). Here, the percentage of small residual vectors decreases to $53.3\%$, indicating that applying a double neural network can be challenging when two different types of operations are involved. A careful design of the Siamese network is therefore necessary. Further analysis is provided in the supplementary document.
VI Discussion
In this work, we have focused on two modern, popular, and off-the-shelf deepfake generation methods: Faceswap-GAN and DiffSwap. Our approach applies a deepfake operation specified by a forensic analyst and uses the norms of resulting residual vectors as a proxy to determine whether the deepfake operation is being applied for the first time or a second time. We examined scenarios where the forensic deepfake operation matches the original deepfake generation method and where the forensic operation differs. Our results reveal that a Siamese detector trained under ideal conditions, where both operations are the same, is also effective when both operations switch to a new type. However, when the two operations differ from each other, the trained Siamese detector is less effective. This suggests the need for a more advanced Siamese detector capable of leveraging processing traces when the two operations are different.
@@ VI Discussion
Compared to end-to-end CNN-based classifiers, our proposed method targets deepfake detection for individuals, with main applications on public figures. Although our method needs training the reconstruction models, the training can be done in advance for each public figure. For example, a journalist can train the reconstruction models for various candidates before they need to verify videos for reporting tasks. Journalists may also share or collaboratively train detectors within their professional networks. To let the detection system support a new individual, the journalist will need to train a reconstruction operator for that individual and then fine-tune the Siamese network.
VII Conclusion and Future Work
@@
In this work, we have proposed to use the method of double neural network operations and individual conditioning for deepfake detection. The proposed detector can achieve better detection performance than end-to-end CNN-based detectors on our curated dataset of public figures with identity labels. We have found that utilizing identity information can make the deepfake detector more reliable. We have also considered scenarios with mismatched first and second deepfake operations for real-world deepfake detection. Our results indicate that a Siamese detector trained on Faceswap-GAN is effective for diffusion-generated deepfake images, provided the additional deepfake operation is also diffusion-based. However, we identified a limitation of the proposed method when the forensic expert’s deepfake operator differs, requiring the training of a new Siamese architecture for that specific combination. In future work, we aim to address this limitation by developing a generalized Siamese detector for deepfake detection.
Supplemental Material
Appendix A Advantage of Identity-Conditioned Feature Extraction
@@
where $S^{(k)}$ is the set of all images belonging to individual $k$, $S_{\text{auth}}$ and $S_{\text{df}}$ are the sets of all authentic and deepfake images, respectively, and $S_{0}$ and $S_{1}$ are the acceptance region and rejection region partitioned by a decision rule [69].
@@
The Bayes risk [69] for an arbitrary rejection region $S_{1}$ is defined as
@@
Standard hypothesis testing technique [69] allows us to derive from (10d) the optimal decision rule that minimizes the Bayes risk or error probability. One can proceed with the derivation and the decision rule turns out to be separable for each individual $k$ and in the form of the likelihood ratio test, namely,
@@
Appendix B Fine-Grained Performance Analysis Over Identities
@@
Appendix C Experimental Validation of Near-idempotence
Our proposed detection method leverages the near-idempotence property of the deepfake operator. Exact idempotence occurs when an altered image, passed through a deepfake generator, depicts no further changes. In the case of near-idempotence, the second operation would lead to small changes compared to the first operation. Let the residues be defined concerning raw data $f$ as follows:
$e_{0}$ $=R_{\text{recon}}(f)-f,$ (16a)
$e_{1}$ $=R_{\text{recon}}(R_{\text{df}}(f))-R_{\text{df}}(f).$ (16b)
To establish near-idempotence, we require $\|e_{0}\|_{2}\gg\|e_{1}\|_{2}$ for all $f$, $R_{\text{df}}$, and $R_{\text{recon}}$, where $R_{\text{df}}$ represents the deepfake operation and $R_{\text{recon}}$ represents the reconstruction operation. In the upcoming two subsections, the first presents the detailed experimental results, while the second offers a short summary.
C-A Detailed Experimental Results
In this subsection, we present detailed experimental results investigating the near-idempotence property of deepfake generators. Specifically, we focus on two types of deepfake operations: Faceswap-GAN (FG) and diffusion-based (D) methods. Based on the choice and the order of deepfake operations applied to an image, we present our results within three categories as follows.
C-A1 $R_{\mathrm{df}}=R_{\mathrm{recon}}=\mathrm{FG}$
We learned a Faceswap-GAN face reconstructor for each of the identities from a subset of the available faces of that identity in the CelebDF dataset. In TABLE V, we present the CDF values of the norm of the feature vector residuals at three threshold points for the first and second operations.
TABLE V: Vector Shifting Statistics for two Faceswap-GAN Operations.
Dataset
Norm of Feature vector residual $\|e\|$
<5.5, <6.0, <6.5
Norm of Siamese Feature Vector Residual $\|e\|$
<1.25, <1.5, <1.75
First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$ First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$
CelebDF 3.74 %, 8.01 %, 16.81 % 97.12 %, 98.80 %, 99.50 % 10.77 %, 18.59 %, 29.36 % 98.78 %, 99.64 %, 99.92 %
CACD 0.03 %, 0.35 %, 1.94 % 72.29 %, 79.86 %, 86.11 % 3.82 %, 7.36%, 13.57 % 62.33 %, 73.19 %, 81.18 %
For near-idempotence-based deepfake image detection to be effective, it is essential to observe a significant difference in the CDF values between the first and second operations. The CDF values for the second operation should be close to $1$ for validating the near-idempotence property. Our results show that this is true for both CelebDF and CACD datasets, with CelebDF demonstrating superior performance. The PDF plots of feature vector residual norms for two Faceswap-GAN operations are shown in the four subfigures of the first column in Fig. 9.
(a) $\mathrm{FG}(f^{\mathrm{cdf}}),\mathrm{FG}(\mathrm{FG}(f^{\mathrm{cdf}}))$
(b) $\mathrm{D}(f^{\mathrm{cdf}}),\mathrm{D}(\mathrm{D}(f^{\mathrm{cdf}}))$
(c) $\mathrm{FG}(f^{\mathrm{cdf}}),\mathrm{FG}(\mathrm{D}(f^{\mathrm{cdf}}))$
(d) $S(\mathrm{FG}(f^{\mathrm{cdf}})),S(\mathrm{FG}(\mathrm{FG}(f^{\mathrm{cdf}})))$
(e) $S(\mathrm{D}(f^{\mathrm{cdf}})),S(\mathrm{D}(\mathrm{D}(f^{\mathrm{cdf}})))$
(f) $S(\mathrm{FG}(f^{\mathrm{cdf}})),S(\mathrm{FG}(\mathrm{D}(f^{\mathrm{cdf}})))$
(g) $\mathrm{FG}(f^{\mathrm{cacd}}),\mathrm{FG}(\mathrm{FG}(f^{\mathrm{cacd}}))$
(h) $\mathrm{D}(f^{\mathrm{cacd}}),\mathrm{D}(\mathrm{D}(f^{\mathrm{cacd}}))$
(i) $\mathrm{FG}(f^{\mathrm{cacd}}),\mathrm{FG}(\mathrm{D}(f^{\mathrm{cacd}}))$
(j) $S(\mathrm{FG}(f^{\mathrm{cacd}})),S(\mathrm{FG}(\mathrm{FG}(f^{\mathrm{cacd}})))$
(k) $S(\mathrm{D}(f^{\mathrm{cacd}})),S(\mathrm{D}(\mathrm{D}(f^{\mathrm{cacd}})))$
(l) $S(\mathrm{FG}(f^{\mathrm{cacd}})),S(\mathrm{FG}(\mathrm{D}(f^{\mathrm{cacd}})))$
Figure 9: PDF plots of residual vector norms. Each subplot is labeled as “Result of first operation, Result of second operation,” where $\mathrm{CDF},\mathrm{CACD},\mathrm{FG},\mathrm{D},S$ denote CelebDF, CACD, Faceswap-GAN, Diffusion, and Siamese network, respectively.
From TABLE V and Fig. 9, we also observe that the feature Vector residual is biased based on the dataset. Consequently, slightly higher threshold values could be chosen for the CACD dataset to better separate the first and second operations. Since the deepfake operator is trained on a portion of the CelebDF dataset, the deepfake images generated using the CelebDF dataset are expected to be more realistic and, hence, dangerous. However, the separation method on vector residual also works better on the CelebDF dataset, mitigating its vulnerability. TABLE V also demonstrates that the Siamese feature vectors with a reduced length of only $50$ can effectively capture the separation capability compared to the original feature vector of length $1856$.
Fig. 10(a) (a) shows the t-SNE visualization of the feature space for authentic, singly processed, and doubly processed images from the CACD dataset.
Figure 10: (a) t-SNE visualization of the feature space of authentic faces from CACD $f$, faceswapped images generated by FaceSwap-GAN $\mathrm{FG}(f)$, and the doubly processed image generated by the learned reconstruction operator $\mathrm{FG}(\mathrm{FG}(f))$. The sample densities of the first and the second classes are visually separable, while the second and third classes exhibit considerable overlap. (b) $e_{0}=\mathrm{FG}(f)-f$ in the t-SNE domain, (c) $e_{1}=\mathrm{FG}(\mathrm{FG}(f))-\mathrm{FG}(f)$ in the t-SNE domain. Vectors in (b) are significantly larger compared to those in (c).
(b) (c)
(a) $e_{0}=\mathrm{FG}(f)-f$ and (b) $e_{1}=\mathrm{FG}(\mathrm{FG}(f))-\mathrm{FG}(f)$, when is Faceswap-GAN (FG) The visualization reveals significant overlap between the features of singly and doubly processed images, while features from authentic images form a separate distribution. We illustrate the vector residual in the t-SNE feature space concerning a single image, from its authentic state to its singly processed version $e_{0}$ in Fig. 10(a) (b), and from the singly processed version to its doubly processed version $e_{1}$ in Fig. 10(a) (c). These plots highlight the vector residuals in the t-SNE space for the CACD dataset derived from the feature values before applying the Siamese network. This figure shows that the first operation results in significant vector residuals, whereas the second operation leads to minimal residuals for $81.3\%$ of the samples. The t-SNE and residual vector plots for the CelebDF dataset are shown in Fig. 11(a), with $94.9\%$ of the residuals being small in the second operation.
Figure 11: (a) t-SNE visualization of the feature space of authentic faces from CelebDF $f$, singly processed images $\mathrm{FG}(f)$, and doubly processed images $\mathrm{FG}(\mathrm{FG}(f))$ generated by the learned reconstruction operator using Faceswap-GAN. The sample densities of the first and second classes are more clearly distinguishable than those of the second and third classes. (b) $e_{0}=\mathrm{FG}(f)-f$ in the t-SNE domain, (c) $e_{1}=\mathrm{FG}(\mathrm{FG}(f))-\mathrm{FG}(f)$ in the t-SNE domain. Vectors in (b) are significantly larger compared to those in (c).
(b) (c)
This better result arises from the reconstruction operator being trained on a subset of the CelebDF dataset.
C-A2 $R_{\mathrm{df}}=R_{\mathrm{recon}}=\mathrm{D}$
To verify the near-idempotence property for diffusion-based deepfake generators, we applied two consecutive diffusion operations to the authentic images. TABLE VI demonstrates that the diffusion-based methods hold the near-idempotence property.
TABLE VI: Vector Shifting Statistics for two Diffusion Operations.
Dataset
Norm of Feature vector residual $\|e\|$
<7.5, <7.75, <8.0
Norm of Siamese Feature Vector Residual $\|e\|$
<1.5, <1.6, <1.7
First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$ First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$
CelebDF 9.74 %, 13.08 %, 17.41 % 80.22 %, 85.58 %, 89.42 % 25.35 %, 29.07 %, 33.06 % 76.84 %, 80.80 %, 84.20 %
CACD 4.55 %, 7.00 %, 10.44 % 90.73 %, 93.18 %, 95.38 % 4.05 %, 5.36 %, 6.61 % 70.31 %, 74.75 %, 79.44 %
Compared to the results of the Faceswap-GAN shown in TABLE V, diffusion-based operators reveal a larger average norm of the feature vector residual. Thus, slightly higher threshold points were chosen in TABLE VII to better emphasize the differences between the first and second operations in the CDF values. At these selected thresholds, the CDF values for the original feature vector for the first operation range from $4.55\%$ to $17.41\%$, while for the second operation, they range from $80.22\%$ to $95.41\%$. Notably, the smaller CDF values for the first operation and the larger values for the second operation in the CACD dataset, compared to the CelebDF dataset, suggest that diffusion images generated from CACD images are more easily detected when using the original feature vectors. For the Siamese feature vector, the norm of the residual feature vector ranges from 4.05% to 33.06% for the first operation and from 70.31% to 84.20% for the second operation. The CDF values approaching $1$ for the second operation again support the near-idempotence property. The PDF plots of feature vector residual norms are shown in the four subfigures of the second column in Fig. 9.
The t-SNE visualizations of the features and vector residual plots are shown in Fig. 12(a). In this case, the near-idempotence property is demonstrated by $89.1\%$ of samples, which is quite an impressive outcome. The significance of this result lies in the fact that the Siamese network was trained using authentic images and deepfake images generated by Faceswap-GAN. Nevertheless, the results in TABLE VI and Fig. 12(a) show that these features can also work with diffusion-based operations. This indicates that the network is capable of extracting meaningful information from the original feature space, even when applied to a different deepfake generation method.
Figure 12: (a) t-SNE visualization of the feature space of authentic faces from CACD $f$, faceswapped images generated by DiffSwap first time $\mathrm{D}(f)$, and second time with repeated operations $\mathrm{D}(\mathrm{D}(f))$. (b) $e_{0}=\mathrm{D}(f)-f$ in the t-SNE domain, (c) $e_{1}=\mathrm{D}(\mathrm{D}(f))-\mathrm{D}(f)$ in the t-SNE domain.
(b) (c)
C-A3 $R_{\mathrm{df}}=\mathrm{D}\,\,\&\,\,R_{\mathrm{recon}}=\mathrm{FG}$
In this case, the first operation is diffusion, whereas the second one is Faceswap-GAN. The results for this scenario are presented in TABLE VII.
TABLE VII: Vector Shifting Statistics for Diffusion Followed by Faceswap-GAN.
Dataset
Norm of Feature vector residual $\|e\|$
<7.75, <8.0, <8.25
Norm of Siamese Feature Vector Residual $\|e\|$
<2.1, <2.2, <2.3
First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$ First Operation $\|e_{0}\|$ Second Operation $\|e_{1}\|$
CelebDF 12.00 %, 16.46 %, 22.73 % 75.93 %, 80.75 %, 84.93 % 54.04 %, 58.28 %, 62.61 % 73.85 %, 77.16 %, 80.20 %
CACD 6.35 %, 9.20 %, 13.23 % 71.01 %, 75.45 %, 79.83 % 26.35 %, 30.66 %, 35.03 % 71.04 %, 75.49 %, 79.27 %
The results indicate that even when the second operation differs from the first, the image is only marginally altered for the second operation in the feature space. This supports our proposed system, as a forensic expert may not always know the exact deepfake method used to generate a fake image. The PDF plots of feature vector residual norms are shown in the four subfigures of the last column in Fig. 9.
However, the feature vector residual in the Siamese feature vector space as shown in Fig. 13(a) reveals that the Siamese network, trained to distinguish between similar deepfake operators, is less effective when two different types of deepfake operations are involved.
Figure 13: (a) t-SNE visualization of the feature space of authentic faces from CACD $f$, faceswapped images generated by DiffSwap $\mathrm{D}(f)$, and further processed by FaceSwap-GAN $\mathrm{FG}(\mathrm{D}(f))$. (b) $e_{0}=\mathrm{FG}(f)-f$ in the t-SNE domain, (c) $e_{1}=\mathrm{FG}(\mathrm{D}(f))-\mathrm{D}(f)$ in the t-SNE domain.
(b) (c)
This suggests the need for properly training the Siamese network regarding the types of deepfake operations in both the first and second stages, as testing on different combinations leads to suboptimal results.
C-B Summary of the Experimental Results
Fig. 11(a) (a) presents the t-SNE features for the CelebDF dataset, which was used to train reconstruction operators. Notably, the reconstruction operators were trained on a subset of the CelebDF dataset, while the results are presented for the remaining subset. Fig. 11(a) (b) and Fig. 11(a) (c) illustrate the t-SNE vector residuals resulting from deepfake operations. The large vector residuals in Fig. 11(a) (b) indicate a significant change in the image features for the first deepfake operation. On the other hand, minimal vector residuals are observed in Fig. 11(a) (c) shows the results when the deepfake operator is applied a second time to the singly processed images. This behavior reflects near-idempotence, supported by $94.9\%$ of the samples. Fig. 12(a) presents a similar result when both operations are diffusion-based. Once again, in this case, the good samples supporting near-idempotence maintain a high percentage of $89.1\%$. Finally, Fig. 13(a) shows the third case, when the first operation is diffusion-based and the second operation is Faceswap-GAN. The results indicate that the good samples are reduced to $53.3\%$ in this case.